If a person who was in energy balance changed their diet such that 20 g of fat replaced 20 g of carbohydrate, what would be the percentage increase in their daily caloric intake? (Show all calculations.) If no other changes were made to the life style, what would be the consequences?
Answer #1

20 GM of fat will gives 20*9.3 kcal =186 kcal.

As one gm of fat gives 9.3 kcal of energy.

20 of carbohydrates contain 20*4 kcal =80 kcal of energy.

So the deficit of energy will be =186-80 kcal = 106 kcal.

So the percentage of change will be (106*100)/186 = 56.98 %

If no other change is happening then , the person will show weight loss because his glycogen store will be depleted first and then fat will also be depleted to meet the energy deficit.

Thank you